Solve equation.
2[x-(2x+4)+3]=2(x+1)

Answer by rfer(16322)

2[x-2x-4+3]=2x+2
2x-4x-8+6]=2x+2
-2x-2=2x+2
-4x=4
x=4/-4
x=-1